a nurse wishes to add water to 30 ounces of a 10% solution of benzalkonium chloride to dilute it to a 4% solution. How much water must she add?
Answer by ewatrrr(24785)

Hi,
*Note: .10*30oz tells us how much benzalkonium chloride is in that total solution
.
.10*30oz = 3oz (there are 3 oz of benzalkonium chloride in that total solution)
.
Question speaks of making a new 4% solution
Let x be the amount of that "total" solution
.
.04*x = 3oz (the actual amount of benzalkonium chloride stays the same)
.
x = 3oz/.04 = 75oz
As we already have 30z of "total" solution, 45oz of water will need to be added
.
check your answer
.04*75oz = the 3 oz of benzalkonium chloride in the "new total" solution